OpenStudy (anonymous):

Two times the first number is one more than the second number if their sum is -64, what are the numbers?

OpenStudy (anonymous):

2x = y + 1 x + y = -64 2x - y = 1 adding 3x = -63 x = -21 and y = 2(-21) -1 = -43

OpenStudy (anonymous):

ok blossom?

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

i let the first number = x and second number = y if you are happy with the answer plz click on good answer box to close the question